# Long Chain-of-Thought for Reasoning Large Language Models

### v1

@@ Abstract
Recent advancements in reasoning with large language models (RLLMs), such as OpenAI-O1 and DeepSeek-R1, have demonstrated their impressive capabilities in complex domains like mathematics and coding. A central factor in their success lies in the application of long chain-of-thought (Long CoT) characteristics, which enhance reasoning abilities and enable the solution of intricate problems. However, despite these developments, a comprehensive survey on Long CoT is still lacking, limiting our understanding of its distinctions from traditional short chain-of-thought (Short CoT) and complicating ongoing debates on issues like "overthinking" and "test-time scaling." This survey seeks to fill this gap by offering a unified perspective on Long CoT. (1) We first distinguish Long CoT from Short CoT and introduce a novel taxonomy to categorize current reasoning paradigms. (2) Next, we explore the key characteristics of Long CoT: deep reasoning, extensive exploration, and feasible reflection, which enable models to handle more complex tasks and produce more efficient, coherent outcomes compared to the shallower Short CoT. (3) We then investigate key phenomena such as the emergence of Long CoT with these characteristics, including overthinking, and test-time scaling, offering insights into how these processes manifest in practice. (4) Finally, we identify significant research gaps and highlight promising future directions, including the integration of multi-modal reasoning, efficiency improvements, and enhanced knowledge frameworks. By providing a structured overview, this survey aims to inspire future research and further the development of logical reasoning in artificial intelligence [^1] Our logo refers to a cute cartoon image - Snake Puppy..
@@ 1 Introduction
In recent years, the emergence of reasoning large language models (RLLMs) such as OpenAI O1 [208] and DeepSeek R1 [155] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [488, 686, 508, 50, 58, 673, 133, 776], as shown in Figure 1. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [798, 437, 439, 421]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by test-time scaling [299, 520, 364]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [440, 385].
@@ 2.2.3 Feasible Reflection for Long CoT
where $\widetilde{n}_{i+1}$ represents the refined version of the subsequent reasoning node $n_{i+1}$, according to the current logic $\texttt{CoT}_{L}^{i}$, feedback result $\mathcal{F}_{i}$, and previous reasoning node $n_{i}$.
@@ 3.1.1 Long CoT External Behavior Analysis
Research shows that contextual examples improve large models’ generative abilities by guiding the formation of reasoning chains [707, 473, 297, 238, 369]. Wang et al. [543] demonstrate that these examples standardize reasoning chain generation relevant to the answers. In an experiment by Madaan et al. [374], removing problem-specific entities from contextual examples, while retaining only the logical structure, led to similar performance as using complete examples, highlighting the logical structure imitation of Long CoT during inference.
More recently, Stechly et al. [484] and Wang and Zhou [579] have shown that modifying the decoding process or designing specific prompts can activate the Long CoT within pre-trained models. They propose that CoT is embedded during pre-training and requires specific activation [658]. Further, Sadr et al. [455] focus the Long CoT source from the training data, and build on this with the notion of “model attribution”, to specifically identify the training data most influential for specific outputs. Building on this, Guo et al. [155] and Xie et al. [622] investigate using rule-based reinforcement learning to directly activate Long CoT during pre-training, aiming to enhance performance [620]. Furthermore, Gandhi et al. [128] identify four key cognitive behaviors, including verification, backtracking, sub-target setting, and backlinking, which successfully facilitate Long CoT. Qwen [649] inherently demonstrates these behaviors, which can be easily triggered by rule-based reinforcement. In contrast, Llama [113] lacks these capabilities and thus requires example-based reinforcement learning to improve significantly.
@@
Recent research has highlighted the upper bounds and limitations of RLLMs across various reasoning tasks [204, 191, 481, 178]. Specifically, Bi et al. [36] investigate these bounds in code generation, showing that RLLMs struggle with tasks that exceed certain complexity thresholds, especially when imitating Long CoT samples of varying complexity. In the context of upper-bound performance, Merrill and Sabharwal [383] and Li et al. [306] focus on single-step arithmetic tasks, concluding that model performance is constrained by input length. Moreover, Feng et al. [118] proposes a mathematical model indicating that fixed-size models cannot produce accurate numerical answers beyond specific limits. However, increasing the number of reasoning steps improves a model’s ability to solve more complex problems.
@@
Research has highlighted the overthinking phenomenon [73, 227, 404, 96, 251], where performance improves with longer reasoning chains up to a threshold, after which it declines. In contrast, Xie et al. [622] and Ma et al. [370] find no significant correlation between reasoning length and accuracy. To explain this, one line of research suggests that Long CoT strategies, like avoiding “snowball errors” [126]. Alternatively, Chen et al. [64], Wolf et al. [602] highlight a performance drop when the reasoning boundaries are exceeded, providing an explanation for the overthinking phenomenon. This suggests that reasoning length and logical complexity should be kept below a certain boundary [756]. Building on this, Wu et al. [611] mathematically determine the feasible reasoning length for Long CoT. Finally, Chen et al. [66] introduces Faraday’s law of Long CoT, which accurately predicts and controls performance.
@@
To understand inference test-time scaling, we will discuss these two paradigms: (1) Vertical Scaling: Vertical scaling involves increasing the reasoning path length. While this can enhance performance, studies by Jin et al. [227] show that, beyond a certain point, longer reasoning paths can degrade performance due to error accumulation. They suggest an optimal path length that depends on the model’s capabilities and task complexity [12, 463]. Furthermore, Chen et al. [64] and Wu et al. [611] explain that excessive exploration lengths beyond the RLLM’s inherent reasoning boundary lead to performance decay, which guides RLLMs for deeper reasoning capabilities [25]. (2) Parallel Scaling: Parallel scaling involves performing multiple reasoning steps and verifying the results. While it shows promise, Parashar et al. [411] and Wang et al. [584] argue that simply increasing inference time does not guarantee improved performance. Wu et al. [610] show that the computational FLOPs $N$ of inference are correlated with the lower bound of performance error, which scales with $\log N$. Additionally, Chen et al. [66] establish an upper bound for parallel scaling, showing that RLLMs cannot exceed Pass@k verification through various verifiers. They further argue that sampling optimization cannot exceed the model’s internal reasoning limitations, demonstrating that for $N$ samples, accuracy is proportional to $\frac{m}{(k/\log N+b)^{2}}$, where $m$, $n$, and $b$ are model-dependent constants.
@@
As RLLMs evolve, it is crucial to differentiate between process supervision and outcome supervision, two key reinforcement learning approaches for complex reasoning tasks, which is essential [632, 123]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [769, 478]. Additionally, PRM faces the issue of reward hacking [10, 101, 403, 24], where agents exploit flaws in the reward function to produce unintended behaviors [155]. Addressing this to surpass rule-based reward systems has become an important research area [155, 622, 419]. Furthermore, Lampinen et al. [260] and Tan [497] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [216] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors.
@@
Earlier, Guo et al. [155] demonstrated that direct RL using rule-based rewards can trigger the aha moment, fostering natural self-reflection without supervision. Following this, Team [507], Xie et al. [622] replicated this phenomenon. Further, Zhou et al. [782] and Meng et al. [382] further extend this phenomenon to multimodal scenarios. However, Liu et al. [346] argue that the aha moment may not emerge in R1-Zero-like training. Instead, they observe that self-reflection patterns, such as superficial self-reflection (SSR), appear at epoch 0, the stage of base models. In this case, self-reflections do not necessarily lead to correct answers. Upon closer examination of R1-Zero training via RL, they find that the increasing response length results not from self-reflection, but from RL optimizing well-designed rule-based rewards.
@@ 3.1.2 Long CoT Internal Mechanism Analysis
Current RLLMs primarily focus on mathematics and coding but have shown potential for generalization to other knowledge-rich domains, sparking growing interest in the mechanism for integrating domain-specific knowledge into Long CoT [608, 622]. Prystawski et al. [429] suggest that generative models store entity knowledge learned during pre-training independently, with the reasoning process in Long CoT linking this knowledge across entities. Radha and Goktas [444] recently introduced the Probabilistic Mixture Model (PMM), which categorizes model outputs into reasoning, memorization, and guessing. They also propose an Information-Theoretic Consistency (ITC) analysis to quantify the relationship between model confidence and strategy selection. Additionally, Jin et al. [228] define "Concept Depth" as the lowest layers at which complex concepts are understood, demonstrating varying levels of knowledge integration in RLLMs. Ou et al. [402] examine RLLM knowledge internalization through knowledge loop evolution, arguing that new knowledge acquisition is shaped by its connection to existing knowledge, with the loop evolving from formation to optimization and from shallow to deep.
@@ 3.2.3 Benchmarks
  • Complex Mathematics: A central focus in complex mathematics is evaluating benchmarks like GSM8K [95] and MATH [172], which assess basic mathematical problem-solving abilities [788, 775]. Recent additions, such as AIME 2024 [6], AIME 2025 [401], MATH-500 [319], AMC 2023 [7], and OlympiadBench [165], expand the evaluation of LLM performance in mathematics. Moreover, Putnam-AXIOM [152] and FrontierMath [140] introduce more complex problems that challenge future reasoning systems. Additionally, ThinkBench [197] and MATH-Perturb [195] focus on robust evaluation for Long CoT.
@@
  • Commonsense Puzzle: Commonsense puzzle benchmarks, including BIG-Bench Hard [495] and ZebraLogic [320], assess models’ ability to reason about commonsense situations. The ARC [87] is often viewed as a challenging commonsense-based AGI test. JustLogic [63] further contributes to the evaluation of deductive reasoning and commonsense problem-solving.
The second focus area concerns Knowledge Benchmarks, essential for evaluating a model’s capability in complex reasoning across various domains:
@@
  • Medical Reasoning: In the realm of Medical Reasoning, the need for complex, domain-specific, and accurate reasoning is paramount [764, 719, 637]. Benchmarks, such as MedQA[225], JAMA Clinical Challenge [55], and Medbullets [55], simulate diagnostic and treatment decision-making processes, reflecting real-world medical practice. These benchmarks evaluate a model’s handling of medical knowledge and reasoning, from diagnosis to treatment planning. Additionally, MedXpertQA [801] introduces a comprehensive evaluation framework combining text and multimodal data, specifically assessing AI’s reasoning capabilities in healthcare.
@@ 3.2.4 Process Evaluations
Recent progress in RLLMs underscores the need for specialized benchmarks to evaluate their deep reasoning abilities in Long CoT [266]. Notably, Lin et al. [320] introduces ZebraLogic, a framework for assessing logical reasoning, especially in complex non-monotonic scenarios. Similarly, BigGSM [64] and GSM-Ranges [472] focus on perturbing numerical values to test logical and arithmetic reasoning in edge cases beyond the models’ training distribution. ROSCOE [142], ReCEval [426], and DiVeRSe [304] are designed to assess each step in the deep reasoning process during Long CoT tasks.
@@ 3.2.5 Advanced Evaluation
Agentic and Embodied reasoning requires models to demonstrate an understanding of real-world interactions, tool use, and adaptive reasoning in response to change. To assess real-world understanding, Wang et al. [568] introduce a benchmark that evaluates agents’ ability to reason about physical concepts. Zhang et al. [745] extend this by assessing agents’ interactions with real-world physics. Additionally, realistic tasks often demand complex planning and tool usage, necessitating benchmarks to evaluate agent reasoning. These benchmarks assess agents’ abilities to navigate and complete tasks in digital environments. Building on this, Huang et al. [191] propose a framework for evaluating decision-making in multi-agent, competitive settings. Nath et al. [393] introduce ToolComp, a benchmark designed to evaluate multi-step tool-use reasoning. To analyze adaptive reasoning in the face of real-world change, OSWorld [623], CogAgent [177], Mobile-Agent-E [589], WebShop [667], WebArena [789], and WebGames [522] assess AI systems across domains such as operating systems, mobile GUIs, browser tasks, and interactive entertainment [770, 559]. Hu et al. [186] present Text2World, which evaluates agents’ ability to generate interactive environments from text to test agent adaptability [695].
@@
Multimodal reasoning refers to a system’s ability to integrate and reason across diverse input types, including text, images, and occasionally code or graphs. This capability is crucial for solving complex problems that require information from diverse formats.
  • Complex Mathematics: Mathematical reasoning often integrates both textual and visual components, such as equations, graphs, or diagrams. Specifically, challenges like MathVista [352], MathVision [561], MathVerse [739], M3CoT-Math [65], CMMaTH [309], EnigmaEval [546], CoMT-Geometry [85], and PGPS9K [737] aim to advance multimodal reasoning in mathematics, improving the evaluation of multimodal Long CoT logic.
  • Complex Code: The second area of focus involves code-related reasoning, where systems interpret textual descriptions and code snippets. Benchmarks like HumanEval-V [728], Code-Vision [550], and Plot2Code [603] evaluate systems’ capabilities to generate or interpret code from natural language and multimodal inputs for assessing systems that integrate natural language processing with programming tasks.
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [351] and M3CoT-Science [65] evaluate how well models combine science information with Long CoT reasoning across various scientific domains.
  • Commonsense Puzzle: This area focuses on commonsense reasoning, where systems combine reasoning cues and images to make deeper conclusions. Chen et al. [65] introduce M3CoT-Commensense, which incorporates commonsense Long CoT reasoning for complex multimodal interactions. Additionally, Wang et al. [544] propose two benchmarks: Clue-Visual Question Answering (CVQA), which tests visual comprehension through three task types, and Clue of Password-Visual Question Answering (CPVQA), which features two task types focusing on the interpretation and application of visual data.

### v3

@@ Abstract
Recent advancements in reasoning with large language models (RLLMs), such as OpenAI-O1 and DeepSeek-R1, have demonstrated their impressive capabilities in complex domains like mathematics and coding. A central factor in their success lies in the application of long chain-of-thought (Long CoT) characteristics, which enhance reasoning abilities and enable the solution of intricate problems. However, despite these developments, a comprehensive survey on Long CoT is still lacking, limiting our understanding of its distinctions from traditional short chain-of-thought (Short CoT) and complicating ongoing debates on issues like "overthinking" and "test-time scaling." This survey seeks to fill this gap by offering a unified perspective on Long CoT. (1) We first distinguish Long CoT from Short CoT and introduce a novel taxonomy to categorize current reasoning paradigms. (2) Next, we explore the key characteristics of Long CoT: deep reasoning, extensive exploration, and feasible reflection, which enable models to handle more complex tasks and produce more efficient, coherent outcomes compared to the shallower Short CoT. (3) We then investigate key phenomena such as the emergence of Long CoT with these characteristics, including overthinking, and test-time scaling, offering insights into how these processes manifest in practice. (4) Finally, we identify significant research gaps and highlight promising future directions, including the integration of multi-modal reasoning, efficiency improvements, and enhanced knowledge frameworks. By providing a structured overview, this survey aims to inspire future research and further the development of logical reasoning in artificial intelligence [^1] Our logo refers to a cute cartoon image - Snake Puppy. Header Image is inspired by Yaoting et al. [759].
@@ 1 Introduction
In recent years, the emergence of reasoning large language models (RLLMs) such as OpenAI O1 [240] and DeepSeek R1 [179] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [549, 777, 572, 60, 68, 761, 155, 880], even passing Turing Test [263], as shown in Figure 1. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [905, 492, 494, 476]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by test-time scaling [338, 584, 413]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [495, 435].
@@ 2.2.3 Feasible Reflection for Long CoT
where $\widetilde{n}_{i+1}$ represents the refined version of the subsequent reasoning node $n_{i+1}$, according to the current logic $\texttt{CoT}_{L}^{i}$, feedback result $\mathcal{F}_{i}$, and previous reasoning node $n_{j}$.
@@ 3.1.1 Long CoT External Behavior Analysis
Research shows that contextual examples improve large models’ generative abilities by guiding the formation of reasoning chains [804, 532, 336, 272, 418]. Wang et al. [609] and Lippmann and Yang [367] demonstrate that these examples standardize reasoning chain generation relevant to the answers both in in-context-learning and supervised-finetuning. In an experiment by Madaan et al. [424], removing problem-specific entities from contextual examples, while retaining only the logical structure, led to similar performance as using complete examples, highlighting the logical structure imitation of Long CoT during inference.
More recently, Stechly et al. [544] and Wang and Zhou [651] have shown that modifying the decoding process or designing specific prompts can activate the Long CoT within pre-trained models. They propose that CoT is embedded during pre-training and requires specific activation [745]. Further, Sadr et al. [512] focus the Long CoT source from the training data, and build on this with the notion of “model attribution”, to specifically identify the training data most influential for specific outputs. Building on this, Guo et al. [179] and Xie et al. [701] investigate using rule-based reinforcement learning to directly activate Long CoT during pre-training, aiming to enhance performance [697]. Furthermore, Gandhi et al. [150] identify four key cognitive behaviors, including verification, backtracking, sub-target setting, and backlinking, which successfully facilitate Long CoT. Qwen [733] inherently demonstrates these behaviors, which can be easily triggered by rule-based reinforcement. In contrast, Llama [134] lacks these capabilities and thus requires example-based reinforcement learning to improve significantly. Moreover, Wang et al. [648] identify a pretraining scaling law that explains how increasing calculation size in RLLMs enhances their reasoning capabilities.
@@
Recent research has highlighted the upper bounds and limitations of RLLMs across various reasoning tasks [236, 220, 541, 203, 143]. Specifically, Bi et al. [46] investigate these bounds in code generation, showing that RLLMs struggle with tasks that exceed certain complexity thresholds, especially when imitating Long CoT samples of varying complexity. In the context of upper-bound performance, Merrill and Sabharwal [433] and Li et al. [345] focus on single-step arithmetic tasks, concluding that model performance is constrained by input length. Moreover, Feng et al. [139] proposes a mathematical model indicating that fixed-size models cannot produce accurate numerical answers beyond specific limits. However, increasing the number of reasoning steps improves a model’s ability to solve more complex problems.
@@
Research has highlighted the overthinking phenomenon [87, 260, 454, 114, 285], where performance improves with longer reasoning chains up to a threshold, after which it declines. In contrast, Xie et al. [701] and Ma et al. [420] find no significant correlation between reasoning length and accuracy. To explain this, one line of research suggests that Long CoT strategies, like avoiding “snowball errors” [148]. Alternatively, Chen et al. [78], Wolf et al. [677] highlight a performance drop when the reasoning boundaries are exceeded, providing an explanation for the overthinking phenomenon. This suggests that reasoning length and logical complexity should be kept below a certain boundary [857]. Building on this, Wu et al. [687] mathematically determine the feasible reasoning length for Long CoT. Finally, Chen et al. [80] introduces Ohm’s law of Long CoT, which accurately predicts and controls performance.
@@
To understand inference test-time scaling, we will discuss these two paradigms: (1) Vertical Scaling: Vertical scaling involves increasing the reasoning path length. While this can enhance performance, studies by Jin et al. [260] show that, beyond a certain point, longer reasoning paths can degrade performance due to error accumulation. They suggest an optimal path length that depends on the model’s capabilities and task complexity [13, 520, 28]. Furthermore, Chen et al. [78] and Wu et al. [687] explain that excessive exploration lengths beyond the RLLM’s inherent reasoning boundary lead to performance decay, which guides RLLMs for deeper reasoning capabilities [29]. (2) Parallel Scaling: Parallel scaling involves performing multiple reasoning steps and verifying the results. While it shows promise, Parashar et al. [462] and Wang et al. [656] argue that simply increasing inference time does not guarantee improved performance. Wu et al. [686] show that the computational FLOPs $N$ of inference are correlated with the lower bound of performance error, which scales with $\log N$. Additionally, Chen et al. [80] establish an upper bound for parallel scaling, showing that RLLMs cannot exceed Pass@k verification through various verifiers. They further argue that sampling optimization cannot exceed the model’s internal reasoning limitations, demonstrating that for $N$ samples, accuracy is proportional to $\frac{m}{(k/\log N+b)^{2}}$, where $m$, $n$, and $b$ are model-dependent constants.
@@
As RLLMs evolve, it is crucial to differentiate between process supervision and outcome supervision, two key reinforcement learning approaches for complex reasoning tasks, which is essential [712, 145, 841]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [871, 538]. Additionally, PRM faces the issue of reward hacking [11, 121, 453, 27, 322], where agents exploit flaws in the reward function to produce unintended behaviors [179]. Addressing this to surpass rule-based reward systems has become an important research area [179, 701, 472]. Furthermore, Lampinen et al. [295] and Tan [559] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [249] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors.
@@
Earlier, Guo et al. [179] demonstrated that direct RL using rule-based rewards can trigger the aha moment, fostering natural self-reflection without supervision. Following this, Team [571], Xie et al. [701] replicated this phenomenon. Further, Zhou et al. [886] and Meng et al. [432] further extend this phenomenon to multimodal scenarios. However, Liu et al. [392] argue that the aha moment may not emerge in R1-Zero-like training. Instead, they observe that self-reflection patterns, such as superficial self-reflection (SSR), appear at epoch 0, the stage of base models. In this case, self-reflections do not necessarily lead to correct answers. Upon closer examination of R1-Zero training via RL, they find that the increasing response length results not from self-reflection, but from RL optimizing well-designed rule-based rewards. Moreover, Yang et al. [744] demonstrate that the "aha moment" is externally marked by increased use of anthropomorphic language during self-reflection and a dynamic adjustment of uncertainty in response to problem difficulty. This process enables the model to maintain reasoning without succumbing to "Reasoning Collapse." Internally, it is characterized by a clear distinction between anthropomorphic traits and logical reasoning, with anthropomorphic language intensifying as the problem becomes more complex.
@@ 3.1.2 Long CoT Internal Mechanism Analysis
Current RLLMs primarily focus on mathematics and coding but have shown potential for generalization to other knowledge-rich domains, sparking growing interest in the mechanism for integrating domain-specific knowledge into Long CoT [684, 701, 872]. Prystawski et al. [484] suggest that generative models store entity knowledge learned during pre-training independently, with the reasoning process in Long CoT linking this knowledge across entities. Radha and Goktas [500] recently introduced the Probabilistic Mixture Model (PMM), which categorizes model outputs into reasoning, memorization, and guessing. They also propose an Information-Theoretic Consistency (ITC) analysis to quantify the relationship between model confidence and strategy selection. Additionally, Jin et al. [261] define "Concept Depth" as the lowest layers at which complex concepts are understood, demonstrating varying levels of knowledge integration in RLLMs. Ou et al. [452] examine RLLM knowledge internalization through knowledge loop evolution, arguing that new knowledge acquisition is shaped by its connection to existing knowledge, with the loop evolving from formation to optimization and from shallow to deep.
@@ 3.2.3 Benchmarks
  • Complex Mathematics: A central focus in complex mathematics is evaluating benchmarks like GSM8K [113] and MATH [197], which assess basic mathematical problem-solving abilities [892, 879]. Recent additions, such as AIME 2024 [7], AIME 2025 [451], MATH-500 [358], AMC 2023 [8], USAMO [475], OlympiadBench [190], and OlympiadArena [231], expand the evaluation of LLM performance in mathematics. Moreover, Putnam-AXIOM [176] and FrontierMath [164] introduce more complex problems that challenge future reasoning systems. Additionally, ThinkBench [226] and MATH-Perturb [224] focus on robust evaluation for Long CoT [33, 782].
@@
  • Commonsense Puzzle: Commonsense puzzle benchmarks, including LiveBench [676], BIG-Bench Hard [556] and ZebraLogic [359], assess models’ ability to reason about commonsense situations. The ARC [105] is often viewed as a challenging commonsense-based AGI test. JustLogic [75] further contributes to the evaluation of deductive reasoning and commonsense problem-solving. Moreover, Li et al. [307] introduce QuestBench, a benchmark designed to evaluate the ability of RLLMs to generate insightful and meaningful questions.
The second focus area concerns Knowledge Benchmarks, essential for evaluating a model’s capability in complex reasoning across various tasks for out of distribution evaluation [623]:
@@
  • Medical Reasoning: In the realm of Medical Reasoning, the need for complex, domain-specific, and accurate reasoning is paramount [866, 816, 718, 468]. Benchmarks, such as MedQA[258], JAMA Clinical Challenge [65], and Medbullets [65], simulate diagnostic and treatment decision-making processes, reflecting real-world medical practice. These benchmarks evaluate a model’s handling of medical knowledge and reasoning, from diagnosis to treatment planning. Additionally, MedXpertQA [908] introduces a comprehensive evaluation framework combining text and multimodal data, specifically assessing AI’s reasoning capabilities in healthcare.
@@ 3.2.4 Process Evaluations
Recent progress in RLLMs underscores the need for specialized benchmarks to evaluate their deep reasoning abilities in Long CoT [301, 897]. Notably, Lin et al. [359] introduces ZebraLogic, a framework for assessing logical reasoning, especially in complex non-monotonic scenarios. Similarly, BigGSM [78] and GSM-Ranges [531] focus on perturbing numerical values to test logical and arithmetic reasoning in edge cases beyond the models’ training distribution. ROSCOE [166], ReCEval [481], and DiVeRSe [343] are designed to assess each step in the deep reasoning process during Long CoT tasks.
@@ 3.2.5 Advanced Evaluation
Agentic and Embodied reasoning requires models to demonstrate an understanding of real-world interactions, tool use, and adaptive reasoning in response to change. To assess real-world understanding, Wang et al. [636] introduce a benchmark that evaluates agents’ ability to reason about physical concepts. Zhang et al. [845] extend this by assessing agents’ interactions with real-world physics. Additionally, realistic tasks often demand complex planning and tool usage, necessitating benchmarks to evaluate agent reasoning. These benchmarks assess agents’ abilities to navigate and complete tasks in digital environments. Building on this, Huang et al. [220] propose a framework for evaluating decision-making in multi-agent, competitive settings. Nath et al. [443] introduce ToolComp, a benchmark designed to evaluate multi-step tool-use reasoning. To analyze adaptive reasoning in the face of real-world change, OSWorld [702], CogAgent [202], Mobile-Agent-E [662], WebShop [754], WebArena [893], and WebGames [586] assess AI systems across domains such as operating systems, mobile GUIs, browser tasks, and interactive entertainment [873, 626, 403]. Hu et al. [213] present Text2World, which evaluates agents’ ability to generate interactive environments from text to test agent adaptability [789].
@@
Multimodal reasoning refers to a system’s ability to integrate and reason across diverse input types, including text, images [248]. This capability is crucial for solving complex problems that require information from diverse formats.
  • Complex Mathematics: Mathematical reasoning often integrates both textual and visual components, such as equations, graphs, or diagrams [728]. Specifically, challenges like MathVista [400], MathVision [628], MathVerse [838], M3CoT-Math [79], CMMaTH [348], EnigmaEval [612], CoMT-Geometry [102], and PGPS9K [835] aim to advance multimodal reasoning in mathematics, improving the evaluation of multimodal Long CoT logic.
  • Complex Code: The second area of focus involves code-related reasoning, where systems interpret textual descriptions and code snippets. Benchmarks like HumanEval-V [825], Code-Vision [616], Plot2Code [678], and ChartMimic [737] evaluate systems’ capabilities to generate or interpret code from natural language and multimodal inputs for assessing systems that integrate natural language processing with programming tasks.
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [399] and M3CoT-Science [79] evaluate how well models combine science information with Long CoT reasoning across various scientific domains [765]. Further, Guo et al. [181] propose MolPuzzle for the evaluation of molecular structure elucidation.
  • Commonsense Puzzle: This area focuses on commonsense reasoning, where systems combine reasoning cues and images to make deeper conclusions. Chen et al. [79] introduce M3CoT-Commensense, which incorporates commonsense Long CoT reasoning for complex multimodal interactions. Further, PuzzleVQA [103] and LEGO-Puzzles [562] focus more on abstract and spatial puzzle reasoning, respectively. Additionally, Wang et al. [610] propose two benchmarks: Clue-Visual Question Answering (CVQA), which tests visual comprehension through three task types, and Clue of Password-Visual Question Answering (CPVQA), which features two task types focusing on the interpretation and application of visual data.